# Current address: ]Susi-Nicoletti-Weg 9, Vienna A-1100, Austria

### v1

@@ Abstract
The phenomenon of induced coherence without induced emission allows to reconstruct the quantum state of a photon that remains undetected. The state information is transferred to its partner photon via optical coherence. Using this phenomenon, a number of established quantum information protocols could be adapted for undetected photons. Despite partial attempts, no general procedure for such adaptation exists. Here we shed light on the matter by showing the close relation between two very dissimilar techniques, namely the quantum state tomography of qubits and the recently developed quantum state tomography of undetected photons. We do so by introducing a set of parameters that quantify the coherence and that mimic the Stokes parameters known from the polarization state tomography. We also perform a thorough analysis of the environment of undetected photons and its role in the reconstruction process.
@@ I Introduction
Recent years have witnessed the advent of quantum technologies fueled by quantum information science that uses intrinsic quantum properties of systems such as photons [1]. These particles are excellent information carriers and can be transmitted through different media, however, they are destroyed in the measurement process. The development of optical coherence has inspired alternative ways to inspect the state of photons without destroying them through a quantum interference effect called induced coherence without induced emission [2, 3]. Among these techniques, one can find the quantification of two-photon transverse momenta [4], entanglement certification of a Bell state [5], and the qubit quantum state tomography [6]. The same interference effect has also inspired several applications for probing objects with undetected light [7], such as imaging [8, 9], spectroscopy [10], optical coherence tomography [11, 12], holography [13], and imaging distillation [14]. Recently, the estimation of the quantum state of undetected photons has been addressed [15].
@@
Even though many of these techniques can be counted among the quantum information protocols, a universal recipe that would allow the translation of already established quantum protocols to the context of coherence-based configuration with undetected photons is missing. In this work, we make a significant step towards filling this gap and delineate the relation between the standard quantum state tomography [16] and the tomography of undetected photons [6]. More specifically, we focus on the quantum state tomography of undetected qubits represented by the polarization of single photons. We reformulate this technique using a novel type of coherence-based parameters that we term visibility Stokes parameters. These parameters are associated with the corresponding visibility operators that resemble Pauli matrices. This way, we make a bridge between the tomography of undetected photons and the standard tomography of qubits. This mutual relation represents a stepping stone for adapting other quantum information techniques to the context of coherence-based quantum operations acting on undetected photons. As a part of our investigation, we present a general analysis of the coherence conditions of the state tomography setup. Unlike in Ref. [6], our treatment does not rely on any extra assumption. [^1] In Ref. [6] an implicit assumption was made that the horizontal polarization mode is perfectly coherent between the two SPDC sources, see also the setup in Fig. 2. This assumption was enforced by the pre-calibration of the experimental setup. To study the effect of the mixedness of the quantum state on the visibility Stokes parameters, we present a comprehensive analysis of the role of the environment.
@@
This manuscript is organized as follows: In Sec. II we review the standard tomography technique with Stokes parameters, Bloch vectors and Pauli operators. Then, in Sec. III, we introduce the visibility Stokes parameters that allow us to represent the pure polarization state of the idler photon by visibility measurements performed on the signal photon. The notion of visibility Stokes parameters is generalized for mixed states in Sec. IV, where a number of complications is discussed and possible remedies are proposed. In Sec. V, we present the quantum operators corresponding to the visibility Stokes parameters, and in Sec. VI, we calculate the post-measurement states of the idler photon. We conclude in Sec. VII.
@@ II Preliminaries
II.1 Quantum state tomography
The method that allows to determine the quantum state $\hat{\rho}_{I}$ of a quantum system such as a photon is called quantum state tomography [16, 17]. In the standard quantum tomography, one has an ensemble of identical systems in state $\hat{\rho}_{I}$ and subjects each system to a measurement on a certain basis, cf. Fig. 1(a). The measurement results provide sufficient information to reconstruct the state $\hat{\rho}_{I}$. For qubits, the measurement bases are usually given as the eigenbases of Pauli operators, and the state can be visually represented in the Bloch sphere depicted in Fig. 1(b), see also next section. If the ensemble is composed of photons, the measurement typically leads to the photons’ destruction.
The original proposal in Ref. [6] reconstructs the quantum state of the idler photon from the visibilities and the phase shift of the interference patterns recorded in succession for the horizontal and vertical polarization modes of the signal photon. Here, we demonstrate, as a side-effect of our investigation, how to record interference patterns in both orthogonal modes simultaneously. Moreover, we adapt the technique in such a way that instead of the visibility and the phase shift in a fixed basis, we measure only the visibilities in the three eigenbases of the Pauli operators. This approach is reminiscent of the standard quantum state tomography.
II.2 Stokes parameters
@@ III Pure states
We begin our discussion by considering pure quantum states of the idler photon. In what follows, we present the method of quantum state tomography of undetected photons that expands the one in Ref. [6]. Most notably, we introduce visibility Stokes parameters that represent a convenient tool for analyzing quantum measurements based on coherence. These new parameters are determined by the visibilities of the recorded interference patterns and thus complement the standard Stokes parameters based on intensities.
The physical setup of the tomography of undetected photons [6] is formed by the imbalanced Zou-Wang-Mandel (ZWM) interferometer. This interferometer, depicted in Fig. 2, consists of two sources of photon pairs, Q1 and Q2. The source Q1 produces photon pairs in a separable state $|\psi_{S}\rangle\otimes|\psi_{I}\rangle$, where $|\psi_{I}\rangle$ is the polarization state of the idler photon that is unknown to us and that we wish to reconstruct
@@
From this expression, we can obtain the analytic formulas for signal intensities and corresponding signal visibilities $\mathcal{V}$ defined as
@@
where $\mathcal{I}_{\textup{max}}$ ($\mathcal{I}_{\textup{min}}$) stands for the intensity maximum (minimum) of the signal photons. The visibilities in all three mutually unbiased bases $H/V$, $D/A$, and $L/R$ turn out to be
@@
Visibilities in bases $D/A$ and $L/R$ depend on the choice of relative phase $\zeta$ in the signal photon’s state in Eq. (16). There does not exist $\zeta$ such that the prefactors in expressions (20) and (21) are both equal, but we can set $\zeta=\pi/2$ and $\zeta=0$ for the $D/A$ and $L/R$ bases, respectively, to obtain
@@
In Appendix A the general formulas for the visibilities in the three bases are shown for arbitrary pumping powers, transmission coefficients, and signal photon states. From Eqs. (19), (22), and (23) it is straightforward to check that
@@
Note that for these visibility Stokes parameters $\mathcal{S}_{x}$, $\mathcal{S}_{y}$, and $\mathcal{S}_{z}$ we use a calligraphic font to differentiate them from the standard Stokes parameters $\textup{S}_{x}$, $\textup{S}_{y}$, and $\textup{S}_{z}$. This distinction will prove important later on when discussing the mixed states. The three visibility Stokes parameters allow us to calculate the coefficients $\alpha$, $\beta$, and the phase $\xi$ of the unknown state in Eq. (15) purely from visibility measurements in different bases. For the case of pure idler states $|\psi_{I}\rangle$, the three parameters $\mathcal{S}_{x}$, $\mathcal{S}_{y}$, and $\mathcal{S}_{z}$ are equal to the three coordinates $(x,y,z)=\vec{r}_{I}$ of the Bloch vector associated with the idler photon’s state. We thus recover the Bloch representation of the idler’s polarization state in Eq. (10). This demonstrates the correspondence between the traditional direct polarization measurement and our visibility measurement of undetected photons.
@@ IV Mixed states
So far, we have considered only pure polarization states of the idler photon. For our tomography technique to be applicable in a practical setting, we present in this section its generalization for mixed states. At first, we present a derivation of the detection probabilities in different measurement bases of the signal photon when the idler photon is in a mixed state. Based on that, we observe that the imbalanced ZWM interferometer in Fig. 2 allows for the identification of both pure and mixed states only with some prior knowledge of the environment. When the environmental conditions are completely unknown, the uncertainty in the reconstruction of mixed states increases with the decoherence in the systems. Then, we derive the visibilities of the obtained interference patterns and define visibility Stokes parameters for mixed states with the aim of being as close as possible to the standard Stokes parameters. Finally, we discuss what assumptions about the setup can be made to enable the mixed state reconstruction in the imbalanced ZWM interferometer.
@@ IV.1 Modelling mixed states with pure states
Mixed states can be represented as pure states living in a larger Hilbert space using a mathematical trick known as purification [1], where a given mixed state $\rho=\sum_{j}p_{j}|\varphi_{j}\rangle\langle\varphi_{j}|$ is associated with a pure state $|\psi\rangle=\sum_{j}\sqrt{p_{j}}|\varphi_{j}\rangle|a_{j}\rangle$, while $\{|a_{j}\rangle\}_{j}$ form an orthonormal basis of an extra ancillary Hilbert space. The partial trace over this extra space then gives the original mixed state as $\rho=\mathrm{Tr}_{A}(|\psi\rangle\langle\psi|)$. The choice of $\{|a_{j}\rangle\}_{j}$ is arbitrary, and any orthonormal basis will do. In the following, we adopt a more physics-inspired approach that bears some similarity to the purification yet shows important differences.
Let us assume without loss of generality that the mixed state $\hat{\rho}_{I}$ of the idler photon results from some decoherence processes in the state-preparation stage, which processes act unitarily on the composite pure state $(\alpha|H\rangle+\beta|V\rangle)\otimes|e\rangle$. Ket $|e\rangle$ denotes the pure state of all the other degrees of freedom of a photon, which we from now on collectively refer to as the environment. The decoherence processes might be polarization-dependent, and so the final state of the idler photon just before entering Q2 attains the form
@@
where $|e_{H}\rangle$ and $|e_{V}\rangle$ are the final states of the environment for the $|H\rangle$ and $|V\rangle$ idler photon polarization mode, respectively. These states are not in general orthogonal, and without loss of generality, we can assume that $0\leq\langle e_{H}|e_{V}\rangle\leq 1$. If we make the following identification
@@
In what follows, we present on a more abstract level the generalization of calculations done for pure states in Sec. III. The results form the stepping stone for discussing mixed state reconstruction, visibilities, and Stokes parameters later on. To account for losses in the setup, the imperfect transmission of objects encountered by the idler photon is modeled as a beam splitter in path $b^{\prime}$ with the reflection coefficient equal to $\sqrt{1-T^{2}}$ that reflects the idler photon into path $w$ [^2] Strictly speaking, $T$ can be polarization-dependent as discussed in Ref. [6]. However, we can also consider the case of homogeneous losses for every polarization component The corresponding term we denote by $|\psi_{w}\rangle$. By $|\psi_{S}\rangle$ we denote the state of the signal photon created in Q1 and propagating along path $a$ after it traversed HWPS and QWPS, i.e., just before it enters the beam splitter BS. Similarly, by $|\psi_{I}\rangle$ we denote the state of the idler photon created in Q1 and propagating along path $b^{\prime}$ after it traversed HWPI and QWPI, which is now of the form of Eq. (31). This state propagates through the source Q2 unaffected and is reflected by the dichroic mirror DMI out of the setup. The state of signal and idler photons just before the projective measurement reads

### v2

# Visibility Stokes parameters as a foundation for quantum information science with undetected photons

@@ Abstract
The framework of measurement operators plays a fundamental role in extracting information about quantum systems. Recently, techniques based on induced coherence have been developed to access the same information for undetected photons. However, there has been a lack of consistent reformulation of quantum operators for these techniques. In this work, we introduce a set of parameters that quantify the polarization of undetected photons based on measured visibilities. Given their similarity to classical counterparts, we refer to them as visibility Stokes parameters. We apply these parameters and the corresponding quantum operators to the problem of quantum state tomography, thoroughly analyzing the environment of undetected photons and its role in the reconstruction process. Because these parameters provide a more intuitive and consistent understanding of the measurement process, we believe that some established quantum information protocols could be adapted for undetected photons.
@@ I Introduction
Recent years have witnessed the advent of quantum technologies fueled by quantum information science that uses intrinsic quantum properties of systems such as photons [1]. These particles are excellent information carriers and can be transmitted through different media, however, they are absorbed in the measurement process. The development of optical coherence has inspired alternative ways to inspect the state of photons without detecting them through a quantum interference effect called induced coherence without induced emission [2, 3]. Among these techniques [4], one can find a quantum eraser [5, 6], the quantification of two-photon transverse momenta [7], entanglement certification of a Bell state [8], and the qubit quantum state tomography [9]. The same interference effect has also inspired several applications for probing objects with undetected light [10], such as imaging [11, 12], spectroscopy [13], optical coherence tomography [14, 15], holography [16], and imaging distillation [17]. Even though many of these techniques can be counted among the quantum information protocols, a universal recipe that would allow the translation of already established quantum protocols to the context of coherence-based configuration with undetected photons is missing.
@@
In this work, we make a step towards filling this gap by delineating the relation between the standard quantum state tomography [18] and the tomography of undetected photons [9]. More specifically, we focus on the quantum state tomography of undetected qubits represented by the polarization of single photons. We reformulate this technique using a novel type of coherence-based parameters that we term visibility Stokes parameters due to their similarity with the Stokes parameters [19]. But while the latter are based on intensity measurements, the former are extracted from visibilities. The resulting quantum state representation by means of visibility Stokes parameters is strikingly similar to the Bloch vector representation of quantum states in standard quantum tomography. This mutual relation between the two tomography techniques represents a stepping stone for adapting other quantum information techniques to the context of coherence-based quantum operations acting on undetected photons. As a part of our investigation, we present a general analysis of the coherence conditions of the state tomography setup. Moreover, we find how coherence and environment are intertwined when mixed states are considered.
@@
This manuscript is organized as follows: In Sec. II we review the standard tomography technique with Stokes parameters, Bloch vectors and Pauli operators. Then, in Sec. III, we introduce the visibility Stokes parameters that allow us to represent the pure polarization state of the idler photon by visibility measurements performed on the signal photon. The notion of visibility Stokes parameters is generalized for mixed states in Sec. IV, where a number of complications is discussed and possible remedies are proposed. In Sec. V some algebraic properties of the visibility Stokes parameters and their connection to standard Stokes parameters are discussed. In Sec. VI, we present the quantum operators corresponding to the visibility Stokes parameters, and in Sec. VII, we calculate the post-measurement states of the idler photon. We conclude in Sec. VIII.
@@ II Preliminaries
The polarization of light and ways to quantify it has a long-standing history in optics [20]. One of the first approaches in classical optics was introduced by G. G. Stokes with the so-called Stokes parameters [19]. Its modern treatment includes the contributions of several authors and concepts such as correlation functions and correlation matrices [21]. In the realm of quantum optics, polarization is measured with the help of the Pauli operators [1] and employing quantum state tomography [18]. In the following, we review the Stokes parameters and the quantum state tomography techniques in more detail.
II.1 Stokes parameters
@@
II.2 Quantum state tomography
The method that allows to determine the quantum state $\hat{\rho}_{I}$ of a quantum system such as a photon is called quantum state tomography [18, 23]. In the standard quantum tomography, one has an ensemble of identical systems in state $\hat{\rho}_{I}$ and subjects each system to a measurement in a certain basis, cf. Fig. 1(a). The measurement results provide sufficient information to reconstruct the state $\hat{\rho}_{I}$. For qubits, the measurement bases are usually given as the eigenbases of Pauli operators, and the state can be visually represented in the Bloch sphere depicted in Fig. 1(b). If the ensemble is composed of photons, the measurement typically leads to the photons’ destruction.
For future reference and to emphasize the differences between the present work and Ref. [9], we list below the individual steps that constitute the quantum state tomography of polarization of undetected photons:
  1. 1.
    A non-linear interferometer (for details see Sec. III) with two sources of photons is employed to produce a pair of photons, the signal and the idler. The effect of induced coherence turns the state of these photons into a superposition of a reference state and the state $\hat{\rho}_{I}$ that is to be reconstructed. This step, which is identical for both Ref. [9] and our work, is schematically represented by the two upper bubbles in Fig. 1(c). Unlike in Ref. [9], our approach considers the most general coherence conditions in the experimental setup.
  2. 2.
    An identical ensemble of photon pairs is produced by the interferometer and subjected to several rounds of measurements. In each round, the signal photons’ state is locally adjusted to a particular form and the signal photons are subjected to measurements in a given polarization basis. This step is schematically represented by the two lower bubbles in Fig. 1(c). While in Ref. [9] the measurement consists of two rounds and the signal photons are always measured in $H/V$ basis, in our work the measurement comprises three rounds and in each round the signal photons are measured in one of Pauli eigenbases $H/V$, $D/A$, and $L/R$.
  3. 3.
    For each round, the phase between the two photon sources is varied and the interference pattern in the rate of signal photons is recorded. The idler photons remain undetected and leave the interferometer, while the interference patterns for signal photon detections are post-processed and relevant quantities are extracted. In Ref. [9], the visibilities and the relative phase shift between two interference patterns were used to reconstruct the polarization state $\hat{\rho}_{I}$. In our work, we instead need only the visibilities of the three interference patterns. This way, the dependence of the reconstructed state on the experimental phase backlash is removed.
Apart from the technical improvements listed above, there is a more conceptual and notable difference — we introduce visibility Stokes parameters that represent a convenient tool for analyzing quantum measurements based on coherence. These new parameters are determined by the visibilities of the recorded interference patterns and can be seen as counterparts of standard Stokes parameters that are based on intensities. The visibility Stokes parameters can be visualized in the form of the visibility Bloch vector and we show its connection to the standard Bloch representation of mixed states. In the upcoming sections we describe our technique in detail and perform a thorough theoretical analysis, first for pure states and then for mixed states.
@@ III Pure states
We begin our discussion by considering pure quantum states of the idler photon. The entire tomography procedure can schematically be described as follows:
  1. 1.
    Idler initial state — The setup shown in Fig. 2 is utilized, where the idler photon with the unknown pure state $|\psi_{I}\rangle$ propagates along (red) path $b^{\prime}$.
  2. 2.
    Signal bases and detection — Three rounds of measurements are performed on an ensemble of identically prepared signal photons. For each round, the signal photon’s state is set to a particular form (see blue box in Fig. 2) and the phase $\phi$ in path $a$ is scanned while detection events are recorded by detectors D1 and D2.
  3. 3.
    Idler state reconstruction — The visibility of the resulting interference pattern for each of the three rounds is determined. The visibility Stokes parameters are calculated from these visibilities, which parameters contain enough information to reconstruct the pure state $|\psi_{I}\rangle$.
In the following, we describe each of these three steps in a separate subsection in more detail.
III.1 Measurement setup
The physical setup of our technique is identical to the one used for the tomography of undetected photons in Ref. [9] and is formed by the imbalanced Zou-Wang-Mandel (ZWM) interferometer. This interferometer, depicted in Fig. 2, consists of two sources of photon pairs, Q1 and Q2. The source Q1 produces photon pairs in a separable state $|\psi_{S}\rangle\otimes|\psi_{I}\rangle$, where $|\psi_{I}\rangle$ is the polarization state of the idler photon that is unknown to us and that we wish to reconstruct
@@ III.1 Measurement setup
III.2 Interference measurements
During the actual tomography measurement, the setup is kept fixed except for the interferometric phase $\phi$. The two detectors D1 and D2 thus record single-photon interference patterns, whose forms can be derived from the expression in Eq. (III.1). The visibility $\mathcal{V}$ of the interference pattern is given by
@@ III.2 Interference measurements
where $\mathcal{I}_{\textup{max}}$ ($\mathcal{I}_{\textup{min}}$) stands for the intensity maximum (minimum). The signal photons can be detected in any polarization basis. For a fixed value of the relative phase $\zeta$ in the signal photon’s state in Eq. (16) one can determine the visibilities in bases $H/V$, $D/A$, and $L/R$. These turn out to be
@@
Visibilities in bases $D/A$ and $L/R$ depend on the choice of relative phase $\zeta$. There does not exist $\zeta$ such that the prefactors in expressions (21) and (22) are all equal, but we can set $\zeta=\pi/2$ and $\zeta=0$ for the $D/A$ and $L/R$ bases, respectively, to obtain
@@
In Appendix A the general formulas for the visibilities in the three bases are shown for arbitrary pumping powers, transmission coefficients, and signal photon states.
III.3 Visibility Stokes parameters
In the previous step, the explicit formulas for the visibilities in three bases $H/V$, $D/A$, and $L/R$ were found. From Eqs. (20), (23), and (24) it is straightforward to check that
@@ III.3 Visibility Stokes parameters
Note that for these visibility Stokes parameters $\mathcal{S}_{x}$, $\mathcal{S}_{y}$, and $\mathcal{S}_{z}$ we use a calligraphic font to differentiate them from the standard Stokes parameters $\textup{S}_{x}$, $\textup{S}_{y}$, and $\textup{S}_{z}$. This distinction will prove important later on when discussing the mixed states. The three visibility Stokes parameters allow us to calculate the coefficients $\alpha$, $\beta$, and the phase $\xi$ of the unknown state in Eq. (15) purely from visibility measurements in different bases.
For the case of pure idler states $|\psi_{I}\rangle$, the three parameters $\mathcal{S}_{x}$, $\mathcal{S}_{y}$, and $\mathcal{S}_{z}$ coincide with the three coordinates $(x,y,z)=\vec{r}_{I}$ of the Bloch vector associated with the idler photon’s state. We thus recover the Bloch representation of the idler’s polarization state in Eq. (10). This demonstrates the correspondence between the traditional direct polarization measurement and our visibility measurement of undetected photons.
@@ IV Mixed states
So far, we have considered only pure polarization states of the idler photon. For our tomography technique to be applicable in a practical setting, we present in this section its generalization for mixed states. We again proceed in a stepwise manner. The procedure is essentially identical to that for pure states with the exception that the theoretical treatment is more involved and there are extra caveats that have to be considered:
  1. 1.
    Idler initial state — The idler photon’s state is now a mixed state $\hat{\rho}_{I}$. In Sec. IV.1 we show how we model this state using a larger pure state. The evolution of both signal and idler photons through the measurement setup is studied in Sec. IV.2.
  2. 2.
    Signal bases and detection — The analytical formulas for detection probabilities in different measurement bases of the signal photon are derived in Sec. IV.2.
  3. 3.
    Idler state reconstruction — The form of visibilities is found in Sec. IV.3. The visibility Stokes parameters for mixed states are defined in Sec. IV.4 with the aim of being as close as possible to the standard Stokes parameters.
  4. 4.
    Role of asymmetric coherence — We observe that when the environmental conditions are completely unknown, the uncertainty in the reconstruction of mixed states increases with the decoherence in the systems. We discuss what additional constraints have to be adopted to enable the mixed state reconstruction in Sec. IV.5.
Each of these steps is discussed in detail below.
@@ IV.1 Modelling mixed states with pure states
In order to treat mixed idler states in our tomography setup we need to represent mixed states in the framework of induced coherence. We do so by modelling mixed states with pure states as explained below. Mixed states can be represented as pure states living in a larger Hilbert space using a mathematical trick known as purification [1], where a given mixed state $\hat{\rho}=\sum_{j}p_{j}|\varphi_{j}\rangle\langle\varphi_{j}|$ is associated with a pure state $|\psi\rangle=\sum_{j}\sqrt{p_{j}}|\varphi_{j}\rangle|a_{j}\rangle$, while $\{|a_{j}\rangle\}_{j}$ form an orthonormal basis of an extra ancillary Hilbert space. The partial trace over this extra space then gives the original mixed state as $\hat{\rho}=\mathrm{Tr}_{A}(|\psi\rangle\langle\psi|)$. The choice of $\{|a_{j}\rangle\}_{j}$ is arbitrary, and any orthonormal basis will do. In the following, we adopt a more physics-inspired approach that bears some similarity to the purification yet shows important differences.
Let us assume without loss of generality that the mixed state $\hat{\rho}_{I}$ of the idler photon results from some decoherence processes in the state-preparation stage, which processes act unitarily on the composite pure state $(\alpha_{0}|H\rangle+\beta_{0}|V\rangle)\otimes|e\rangle$. The vector $|e\rangle$ denotes the pure state of all the other degrees of freedom of a photon, which we from now on collectively refer to as the environment. The decoherence processes might be polarization-dependent, and so the final state of the idler photon just before entering Q2 attains the form
@@
where $|e_{H}\rangle$ and $|e_{V}\rangle$ are the final states of the environment for the $|H\rangle$ and $|V\rangle$ idler photon polarization mode, respectively. These states are not in general orthogonal and without loss of generality we can assume that $0\leq\langle e_{H}|e_{V}\rangle\leq 1$. If we make the following identification
@@
We assume that the $H$- and $V$-polarized terms in Bell state $|\Psi\rangle$ are perfectly coherent and so they share the same environmental state $|e_{\Psi}\rangle$.
IV.2 Photon evolution through setup
In what follows, we present on a more abstract level the generalization of calculations done for pure states in Sec. III.1. The results form the stepping stone for discussing mixed state reconstruction, visibilities, and Stokes parameters later on. To account for losses in the setup, the imperfect transmission of objects encountered by the idler photon is modeled as a beam splitter in path $b^{\prime}$ with the reflection coefficient equal to $\sqrt{1-T^{2}}$ that reflects the idler photon into path $w$ [^1] Strictly speaking, $T$ can be polarization-dependent as discussed in Ref. [9]. However, here we consider the case of homogeneous losses for every polarization component. The corresponding term we denote by $|\psi_{w}\rangle$. By $|\psi_{S}\rangle$ we denote the state of the signal photon created in Q1 and propagating along path $a$ after it traversed HWPS and QWPS, i.e., just before it enters the beam splitter BS. Similarly, by $|\psi_{I}\rangle$ we denote the state of the idler photon created in Q1 and propagating along path $b^{\prime}$ after it traversed HWPI and QWPI, which is now of the form of Eq. (32). This state propagates through the source Q2 unaffected and is reflected by the dichroic mirror DMI out of the setup. The state of signal and idler photons just before the projective measurement reads